# Cycles in spherical Deligne complexes and application to $K(\pi,1)$-conjecture for Artin groups

### v1

@@ 11 $K(\pi,1)$ for some higher-dimensional families
Denote the vertices of $\Lambda$ adjacent to $s$ by $\{b_{i}\}_{i=1}^{n}$. Let $e_{i}$ be the edge labeled by $4$ that contains $b_{i}$, whenever such an edge exists. Let $t_{i}$ be the valence one vertex of $e_{i}$. For $1\leq i,j,k\leq n$, let $\Lambda_{ijk}$ be the subgraph of $\Lambda$ spanned by $a,b_{i},b_{j},b_{k}$. Similarly we define subgraphs $\Lambda_{ij}$ and $\Lambda_{ijk\ell}$ of $\Lambda$. Let $\Lambda_{e_{i},j}=\Lambda_{ij}\cup e_{i}$, whenever $e_{i}$ exists. Let $\Lambda_{e_{i},e_{j}}=\Lambda_{ij}\cup e_{i}\cup e_{j}$ whenever $e_{i}$ and $e_{j}$ exist. We claim that
  1. 1.
    for any pairwise distinct $\{i,j,k\}$, the $(i,j)$-subdivision, $(i,k)$-subdivision and $(j,k)$-subdivision of $\Delta_{\Lambda,\Lambda_{ijk}}$ are all bowtie free and downward flag;
  2. 2.
    the vertex set of $\Delta_{\Lambda,\Lambda_{e_{i},j}}$, endowed with the order induced from the linear order in $\Lambda_{e_{i},j}$ with $t_{i}$ being the largest vertex, is bowtie free and upward flag.
Let $m$ be the number of edges in $\Lambda$ labeled by $4$. Note that $m\leq n$. We prove the claim by induction on $n+m$. The base cases are $n=3$ and $m=0$, where Assertion 1 follows from Theorem 10.1 and Assertion 2 is empty; and $n=2$ and $m=1$, where case Assertion 2 follows from Theorem 3.22 and Assertion 1 is empty.
First we show Assertion 2 of the claim. As $n\geq 3$, we can find $k\notin\{i,j\}$. Consider $\Lambda^{\prime}=e_{i}\cup\Lambda_{ijk}$. By our induction assumption and Proposition 3.32, the $(b_{j},b_{k})$-subdivision of $\Delta_{\Lambda,\Lambda^{\prime}}$ satisfies all the assumptions of Theorem 3.27. By Proposition 8.1 (1) and (2), the vertex set of $\Delta_{\Lambda,\Lambda_{e_{i},j}}$ with the desired order is a bowtie free and upward flag poset, and the $(b_{j},b_{k})$-subdivision of $\Delta_{\Lambda,\Lambda_{ijk}}$ is bowtie free and downward flag. By Proposition 8.2 and our induction assumption, the $(b_{i},b_{j})$-subdivision and $(b_{j},b_{k})$-subdivision of $\Delta_{\Lambda,\Lambda_{ijk}}$ are also bowtie free and downward flag.
Now we show Assertion 1 of the claim. If at least one of $\{e_{i},e_{j},e_{k}\}$ exists, then Assertion 1 follows from the same argument in the previous paragraph. Now we assume none of $\{e_{i},e_{j},e_{k}\}$ exists. If $n=3$, then we must have $m=0$, which is one of the base cases of the induction. Now assume $n>3$. Then there is $\ell\notin\{i,j,k\}$. Now we consider $\Delta_{\Lambda,\Lambda_{ijk\ell}}$. By induction, all the assumptions of Proposition 3.34 are satisfied (with $n=1$ in Proposition 3.34). Thus Proposition 8.8 (1) and (2) imply that Assertion 1 holds for a particular subdivision of $\Delta_{\Lambda,\Lambda_{ijk}}$. Other subdivisions can be treated by symmetry.
Let $\mathcal{C}_{2}$ be the class of Dynkin diagrams satisfying the two requirements of Theorem 11.5. Let $\mathcal{C}_{1}=\{\widetilde{B}_{4},\widetilde{C}_{4},\widetilde{D}_{4}\}$. We now verify all the assumptions of [Hua23, Proposition 7.2] hold. Assumption 1 of [Hua23, Proposition 7.2] follows from the above claim, Proposition 3.35, Theorem 3.27, and Proposition 3.34. Assumption 2 of [Hua23, Proposition 7.2] is clear. Let $\Lambda^{\prime}$ be as in Assumption 3 of [Hua23, Proposition 7.2]. Then $\Lambda^{\prime}$ is spherical, hence satisfies $K(\pi,1)$-conjecture by [Del72]. Now we are done by [Hua23, Proposition 7.2]. ∎
Appendix A Proposed metric on the core
We propose two metrics on the complex $\Delta_{S,S^{\prime}}$ discussed in Conjecture 1.13: one conjectured to be CAT$(0)$, and another conjectured to satisfy other notions of non-positive curvature [Lan13, DL15, CCHO14, CCG+20]. The first one is easier to define, but harder to work with (we actually do not work with it); the second one takes more effort to define, but easier to work with.
For an Artin group $A_{S}$, we use $S$ to denote the standard generating set, $\Gamma$ to denote the presentation graph, and $\Lambda$ to denote the Dynkin diagram. We say the Artin group $A_{S}$ dominates another Artin group $A_{S^{\prime}}$ if there exists an isomorphism of their presentation graphs $f:\Gamma\to\Gamma^{\prime}$ such that the label of each edge $e\subset\Gamma$ is $\geq$ the label of $f(e)$. The following is a consequence of [Lan50b].
Lemma A.1.
Suppose $A_{S}$ is an irreducible almost spherical type Artin group. Then
  1. 1.
    either $A_{S}$ dominates an irreducible Artin group of Euclidean type, or $A_{S}$ has Dynkin diagram $[3,5,3]$ or $[5,3,3,3]$;
  2. 2.
Here $[3,5,3]$ is the linear Dynkin diagram whose consecutive edges are labeled by $3,5$ and $3$. Similarly we define $[5,3,3,3]$.
Definition A.2 (Proposed metric I).
Given an Artin group $A_{S}$, and let $S^{\prime}\subset S$ be an irreducible almost spherical subset. To metrize the relative Artin complex, it suffices to metrize the fundamental domain with respect to the action $A_{S}\curvearrowright\Delta_{S,S^{\prime}}$, which is a simplex $K_{S^{\prime}}$ whose vertices are of type $\hat{s}$ with $s\in S^{\prime}$. Now we choose a shape for this fundamental domain as follows.
Let $A_{S^{\prime\prime}}$ be an Artin group dominated by $A_{S^{\prime}}$ under a bijection $f:S^{\prime}\to S^{\prime\prime}$ such that $A_{S^{\prime\prime}}$ is either Euclidean, or $[3,5,3]$, or $[5,3,3,3]$. Then associated Coxeter group $W_{S^{\prime\prime}}$ acts properly and cocompactly by isometries on a metric space $X$ with $X$ being either $\mathbb{E}^{n}$ or $\mathbb{H}^{3}$ or $\mathbb{H}^{4}$. The reflection hyperplanes (i.e. fixed points of conjugates of generators of $W_{S^{\prime\prime}}$) cut $X$ into a simplicial complex, which is isomorphic to the Coxeter complex $\mathsf{C}_{S^{\prime\prime}}$. Let $K_{S^{\prime\prime}}$ be the fundamental domain of the action $W_{S^{\prime\prime}}\curvearrowright\mathsf{C}_{S^{\prime\prime}}$, which is a simplex with an inherit metric from $X$. Now we identify $K_{S^{\prime}}$ with $K_{S^{\prime\prime}}$ with the vertex of type $\hat{s}$ in $K_{S^{\prime}}$ with $s\in S^{\prime}$ is identified with the vertex of type $\hat{f}(s)$ in $K_{S^{\prime\prime}}$. This gives a metric on $K_{S^{\prime}}$, hence induces a piecewise Euclidean or hyperbolic metric in $\Delta_{S,S^{\prime}}$.
Conjecture A.3.
Let $A_{S}$ be an Artin group and suppose $S$ contains a subset $S^{\prime}$ such that $A_{S^{\prime}}$ is almost spherical. Then $\Delta_{S,S^{\prime}}$ with the metric in Definition A.2 is CAT$(0)$ when $S^{\prime\prime}$ in Definition A.2 is Euclidean, and is CAT$(-1)$ when $S^{\prime\prime}$ in Definition A.2 is either of type $[3,5,3]$ or $[5,3,3,3]$. Hence $\Delta_{S,S^{\prime}}$ is contractible.
To prove Conjecture A.3, it suffices to show $\Delta_{S,S^{\prime}}$ is simply-connected, which follows from Lemma 3.4, and the link of each vertex is CAT$(1)$. The link condition is rather challenging to verify, especially when $\dim(\Delta_{S,S^{\prime}})\geq 2$. We refer to [EM04, EM02, McC09] for some related discussion.
Due to the difficulty of verifying the link condition in the CAT$(0)$ setting, we propose an alternative metric on $\Delta_{S,S^{\prime}}$ for most types of almost spherical sets $S^{\prime}$, which is not exactly CAT$(0)$, but has strong flavor of non-positive curvature.
Definition A.4 (Proposed metric II).
Let $S,S^{\prime},\Delta_{S,S^{\prime}},K_{S^{\prime}}$ be as in Definition A.2. We assume the type of $S^{\prime}$ does not belong to $\{\widetilde{F}_{4},\widetilde{E}_{6},\widetilde{E}_{7},\widetilde{E}_{8},[3,5,3],[5,3,3,3]\}$. By Lemma A.1 (2), there is an Artin group $A_{S^{\prime\prime}}$ dominated by $A_{S^{\prime}}$ under a bijection $f:S^{\prime}\to S^{\prime\prime}$ such that the type of $A_{S^{\prime\prime}}$ belongs to $\{\widetilde{G}_{2},\widetilde{A}_{n},\widetilde{B}_{n},\widetilde{C}_{n},\widetilde{D}_{n}\}$. Let $K_{S^{\prime\prime}}$ be as in Definition A.2. By the same procedure as in Definition A.2, it suffices to metrize $K_{S^{\prime\prime}}$.
If $S^{\prime\prime}$ is of type $\widetilde{G}_{2}$, then we metrize in the same we as before. If $S^{\prime\prime}$ is of type $\widetilde{B}_{n},\widetilde{C}_{n},\widetilde{D}_{n}$, then $W_{S^{\prime\prime}}$ acts properly and cocompactly on $\mathbb{E}^{n}$ as before, and this action is isometric with respect to the $\ell^{\infty}$-metric. We equipped $K_{S^{\prime\prime}}$ with the induced $\ell^{\infty}$-metric from $\mathbb{E}^{n}$. If $S^{\prime\prime}$ is of type $\widetilde{A}_{n}$, then the Coxeter complex $\mathsf{C}_{S^{\prime\prime}}$ can be identified with the subset $H=\{x\in\mathbb{E}^{n+1}\mid x_{1}+x_{2}+\cdots+x_{n+1}=0\}$ cut out by the hyperplanes $x_{i}-x_{j}\in\mathbb{Z}$ for $11\leq i\neq j\leq n+$ 1, with a fundamental domain $K_{S^{\prime\prime}}$ be the set $\{x_{1}\leq x_{2}\leq\cdots\leq x_{n}\leq x_{1}+1\}$ in $H$. We equip $\mathbb{E}^{n+1}$ with $\ell^{\infty}$-metric, and metrize $K_{S^{\prime\prime}}$ with the induced metric.
The idea of domination allows us to “pretend” $S^{\prime}$ is of Euclidean type, and metrize with $\ell^{\infty}$-metric, even if it is actually of hyperbolic type. However, as $[3,5,3]$ and $[5,3,3,3]$ do not dominate any Euclidean type, we need new candidates of metric in these two cases. In the special case of $S=S^{\prime}\in\{\widetilde{A}_{n},\widetilde{C}_{n}\}$, the metric in Definition A.4 was defined in [Hae21].
Here is a more concrete description of the $\ell^{\infty}$-metric on $K_{S^{\prime\prime}}$ when $S^{\prime\prime}$ is of type $\{\widetilde{B}_{n},\widetilde{C}_{n},\widetilde{D}_{n}\}$. In the case of type $\widetilde{C}_{n}$, the Coxeter complex $\mathsf{C}_{S^{\prime\prime}}$ is isomorphic to $\mathbb{E}^{n}$ cut out by the hyperplanes $x_{i}\pm x_{j}\in\mathbb{Z}$ and $x_{i}\in\frac{1}{2}\mathbb{Z}$. Then the fundamental domain $K_{S^{\prime\prime}}$ can be chosen to be the subset of $\mathbb{E}^{n}$ defined by $0\leq x_{1}\leq x_{2}\leq\cdots\leq x_{n}\leq\frac{1}{2}$, equipped with the $\ell^{\infty}$ metric. If $S^{\prime\prime}$ is of type $\widetilde{B}_{n}$, then the subdivision of the associated Coxeter complex $\mathsf{C}_{S^{\prime\prime}}$ as in Definition 3.30 is isomorphic to the Coxeter complex of type $\widetilde{C}_{n}$, thus a type $\widetilde{B}_{n}$ fundamental domain is the union of two copies of type $\widetilde{C}_{n}$ fundamental domain along an appropriate codimension 1 face. If $S^{\prime\prime}$ is of type $\widetilde{D}_{n}$, then subdivision of the associated Coxeter complex $\mathsf{C}_{S^{\prime\prime}}$ as in Definition 3.33 is isomorphic to the Coxeter complex of type $\widetilde{C}_{n}$. Hence a $\widetilde{D}_{n}$-fundamental domain is the union of four copies of $\widetilde{C}_{n}$ fundamental domain.
Recall that a geodesic bicombing in a metric space $X$, is the assignment of a geodesic segment from $x$ to $y$, for each order pair of points $(x,y)$ in $X$. We do not require the geodesic segment from $x$ to $y$ is the same as the geodesic segment from $y$ to $x$. Note that in Definition A.4, each top-dimensional simplex has geodesic bicombing coming from the linear structure on each simplex. The hope is that these geodesic bicombing fit together to form a geodesic bicombing on the whole space such that the bicombing varies continuously with respect to their endpoints, hence one can show $\Delta_{S,S^{\prime}}$ is contractible by using the geodesic contraction.
Conjecture A.5.
Let $A_{S}$ be an Artin group and suppose $S$ contains a subset $S^{\prime}$ which is almost spherical, but its type is not contained in $\{\widetilde{F}_{4},\widetilde{E}_{6},\widetilde{E}_{7},\widetilde{E}_{8},[3,5,3],[5,3,3,3]\}$. Then $\Delta_{S,S^{\prime}}$ equipped with the metric in Definition A.4 is a metric space with convex geodesic bicombing in the sense of [DL15], hence contractible.
We conjecture that similar to the situation of proving Conjecture A.3, one is reduced to check an appropriate form of link condition. Again checking this link condition will be the most challenging part. Though we believe the link condition to check in the context of Conjecture A.5 will be much simpler than Conjecture A.3 - when $S^{\prime\prime}$ is not of type $\widetilde{G}_{2}$, then link condition should only involve cycles in the 1-skeleton of the link up to length $6$. Combined with some existing link condition in [Hae22], we know the following.
Lemma A.6.
Suppose $\Lambda$ is a connected Dynkin diagram, and $\Lambda^{\prime}$ is an admissible subgraph of $\Lambda$ in the sense of Definition 3.13. Suppose $\Lambda^{\prime}$ is almost spherical with vertex set $S^{\prime}$ and $\Lambda$ has vertex set $S$. Suppose the type of $S^{\prime}$ is not one of $\{\widetilde{G}_{2},\widetilde{F}_{4},\widetilde{E}_{6},\widetilde{E}_{7},\widetilde{E}_{8},[3,5,3],[5,3,3,3]\}$. Then there is a criterion only involving cycles in the 1-skeleton up to length 6 in the link of each vertex of $\Delta_{S,S^{\prime}}$ such that as long as such link criterion is satisfied, then $\Delta_{S,S^{\prime}}$ is contractible. More precisely,
  1. 1.
    if $S^{\prime}$ dominates type $\widetilde{A}_{n}$ (in which case the Dynkin diagram of $S^{\prime}$ must be a cycle, hence elements in $S^{\prime}$ has a natural cyclic order), and then we need to check $\Delta_{S,S^{\prime}}$, viewed as a simplicial complex of type $S^{\prime}$, satisfies the assumptions of Theorem 3.28;
  2. 2.
    if $S^{\prime}$ dominates type $\widetilde{C}_{n}$ (in which case the Dynkin diagram of $S^{\prime}$ is a line, hence $S^{\prime}$ has two linear orders), then we need to check if $S^{\prime}$ is equipped with one of these linear orders, then $\Delta_{S,S^{\prime}}$ satisfies the assumptions of Theorem 3.27;
  3. 3.
    if $S^{\prime}$ dominates type $\widetilde{B}_{n}$ (in which case the Dynkin diagram of $S^{\prime}$ is isomorphic to Figure 5 left), then we need to check assumptions of Proposition 3.32 are satisfied;
  4. 4.
    if $S^{\prime}$ dominates type $\widetilde{D}_{n}$ (in which case the Dynkin diagram of $S^{\prime}$ is isomorphic to Figure 5 right), then we need to check assumptions of Proposition 3.34 are satisfied.
In the case when $S^{\prime}$ is of type $\widetilde{G}_{2}$, there is an obvious link condition to check coming from the CAT$(0)$ link condition.
Note that Lemma A.6 requires $\Lambda^{\prime}$ to be an admissible subgraph of $\Lambda$. If $\Lambda^{\prime}$ is not an admissible subgraph of $\Lambda$, then the link criterion in Lemma A.6 does not apply directly. For example, in Case 2 of Lemma A.6, if $\Lambda^{\prime}$ is not an admissible, then $\Delta_{S,S^{\prime}}$ with the relation induced from the linear order from $S^{\prime}$ is not a poset. So Theorem 3.27 does not apply directly. Similar problems happen with the other cases. This leads to us to formulate the following somewhat less precise conjecture.
Conjecture A.7.
Under the same assumption of Lemma A.6, even if $\Lambda^{\prime}$ is not admissible in $\Lambda$, there still exists a criterion only involving cycles in the 1-skeleton up to length 6 in the link of each vertex of $\Delta_{S,S^{\prime}}$ such that as long as such link criterion is satisfied, then $\Delta_{S,S^{\prime}}$ is contractible.
For the remaining types of $\Lambda^{\prime}$, we ask the following.
Question A.8.
Suppose $\Lambda$ is a connected Dynkin diagram, and $\Lambda^{\prime}$ is an induced subgraph of $\Lambda$ in the sense of Definition 3.13. Suppose the type of $\Lambda^{\prime}$ is one of $\{\widetilde{F}_{4},\widetilde{E}_{6},\widetilde{E}_{7},\widetilde{E}_{8}$, [3,5,3],[5,3,3,3]}[3,5,3],[5,3,3,3]\}[ 3 , 5 , 3 ] , [ 5 , 3 , 3 , 3 ] }. Is it true that there is a criterion only involving cycles in the 1-skeleton up to a certain length in the link of each vertex of $\Delta_{S,S^{\prime}}$ such that as long as such link criterion is satisfied, then $\Delta_{S,S^{\prime}}$ is contractible?
Appendix B Conjectures on cycles in relative Artin complexes
We list two conjectures which are necessary for the strategy in [Hua23] to work. We refer to Section 3.2 and Section 3.3 for related terminologies.
Conjecture B.1.
Suppose $A_{S}$ is an irreducible spherical Artin group with Dynkin diagram $\Lambda$. Let $\Lambda^{\prime}$ be a linear subdiagram of $\Lambda$ with consecutive vertices $\{s_{i}\}_{i=1}^{n}$ such that the edge between $s_{n-1}$ and $s_{n}$ has label $\geq 4$. Then the vertex set of $\Delta_{\Lambda,\Lambda^{\prime}}$ equipped with the relation induced from $s_{1}<s_{2}<\cdots<s_{n}$ (as in Definition 3.12) is a bowtie free and upward flag poset.
The bowtie free part is already known, and is a consequence of [Hua23, Proposition 2.8]. For the upward flag part, the case when $A_{S}=A_{S^{\prime}}$ is type $B_{n}$ follows from [Hae21]; the case that $A_{S}$ is of type $F_{4}$ and $A_{S^{\prime}}$ is of type $B_{3}$ follows from Proposition 4.2; the case $A_{S}$ being type $H_{3}$ follows from Theorem 7.1. The remaining case of the conjecture is the upward flag part of $A_{S}$ being type $H_{4}$.
Conjecture B.2.
Suppose $A_{S}$ is an irreducible spherical Artin group. Let $S^{\prime}\subset S$ such that $A_{S^{\prime}}$ has Dynkin diagram isomorphic to the type $D_{n}$ Dynkin diagram for $n\geq 3$ (the isomorphism does not need to preserve edge labels). Let $\{b_{i}\}_{i=1}^{n+1}$ be vertices in $S^{\prime}$ as in Figure 5 left. Then the $(b_{1},b_{2})$-subdivision of $\Delta_{S,S^{\prime}}$ (in the sense of Definition 3.30) is a bowtie free and downward flag poset.
Note that we specifically allow the case when the Dynkind diagram $A_{S^{\prime}}$ is isomorphic to the type $D_{3}$ diagram (though edge labels might not be preserved). While $D_{3}$ diagram is the same as $A_{3}$ diagram, but we are considering a subdivision of $\Delta_{S,S^{\prime}}$ by viewing the Dynkin diagram as $D_{3}$ rather than $A_{3}$, as explained in Definition 3.30.
Again the bowtie free part follows from [Hua23, Proposition 2.8]. The downward flag part is known when $n=3$, and $A_{S}$ is of type $A_{n},B_{n},H_{3},F_{4}$ this is a combination of Lemma 11.2, Theorem 5.6, Theorem 7.1 and Proposition 4.2. By Theorem 3.16, the downward flag part is also known when $A_{S}$ is of type $D_{n}$, and $S^{\prime}=\{b_{1},b_{2},b_{3}\}$ in Figure 5. The downward flag part is true when $A_{S}=A_{S^{\prime}}$ is of type $D_{4}$, see Theorem 10.1. The remaining cases are open.
Conjecture B.2 generalizes an earlier conjecture of Haettel, in the special case of $A_{S}=A_{S^{\prime}}$ are both of type $D_{n}$.
Our plan is to first establish Conjecture B.1 and Conjecture B.2. Then use a suitable propagation argument to establish the following.
Conjecture B.3.
Conjecture B.1 and Conjecture B.2 are still true for any $A_{S}$ with connected Dynkin diagram $\Lambda$, and $A_{S^{\prime}}$ corresponds to an admissible full subgraph of $\Lambda$ satisfying the requirements in Conjecture B.1 and Conjecture B.2.
Then one can use Conjecture B.3 together with Lemma A.6 to establish Conjecture A.5 partially. The reason for the word “partially” is that Conjecture B.3, Theorem 3.28, Theorem 3.27, Proposition 3.32 and Proposition 3.34 require the subgraph $\Lambda^{\prime}$ inside $\Lambda$ to be admissible, this will ensure the relation on the vertices of the associated complex to be a poset. Actually, the poset part of Conjecture B.1 and Conjecture B.2 will fail if the subdiagram $\Lambda^{\prime}$ of $\Lambda$ corresponding to $A_{S^{\prime}}$ is not admissible. In this case, we need to find suitable replacements in order to prove Conjecture A.5 fully.

### v3

@@ 12 $K(\pi,1)$ for some higher-dimensional families
The $n\leq 2$ case follows from [CD95]. The $n=3$ case follows from Corollary 10.4. For $n=4$, the only possible Coxeter diagrams are shown in Figure 16. The $K(\pi,1)$-conjecture for diagrams (1), (2) and (4) are proved in [Hua24a]. The remaining diagrams are consequences of Proposition 12.1 (the choice of $\Lambda^{\prime}$ is the thickened subgraphs in Figure 16). ∎
Appendix A Injective metric on type $B_{n}$ Falk complex
In this section, $\Lambda$ will be the Coxeter diagram of type $B_{n}$. Let consecutive nodes in $\Lambda$ be $\{s_{1},\ldots,s_{n}\}$ with $m_{s_{n-1},s_{n}}=4$. Let $\mathcal{A}$ be the reflection arrangement of type $B_{n}$ in $\mathbb{R}^{n}$. Then hyperplanes of $\mathcal{A}$ are of form $\{x_{i}=0\}_{i=1}^{n}$ and $\{x_{i}=x_{j}\}_{1\leq i\neq j\leq n}$. Let $H$ be the hyperplane $x_{1}=0$. Then a simple calculation implies that deconing of $\mathcal{A}$ with respect to $H$ gives the affine arrangement $\mathcal{B}$ in $\mathbb{R}^{n-1}$ made of $y_{i}=-1,0,1$ for $11\leq i\leq n-$ 1, and $y_{i}=\pm y_{j}$ for $11\leq i\neq j\leq n-$ 1.
Let $S_{\mathcal{A}}$ and $D_{\mathcal{B}}$ be defined in Section 4.3. Then $S_{\mathcal{A}}$ is isomorphic to the Coxeter complex of the associated Coxeter group, hence each vertex of $S_{\mathcal{A}}$ has a type $\hat{s}_{i}$ for some $i$. The complex $D_{\mathcal{B}}$ is a subdivision of $[-1,1]^{n-1}$ as follows: first we subdivide into $2^{n-1}$ unit cubes, then we subdivide each of these unit cubes into $(n-2)!$ copies of $n$-dimensional orthoschemes, such that each of the orthoschemes contain $(0,\ldots,0)$. Note that $D_{\mathcal{B}}$ can be realized as the maximal subcomplex of $S_{\mathcal{A}}$ which is contained in the interior of a hemisphere bounded by $H\cap S_{\mathcal{A}}$. Thus it makes sense to talk about types of vertices of $D_{\mathcal{B}}$, using this embedding $D_{\mathcal{B}}\to S_{\mathcal{A}}$. It follows that the vertex $(0,\ldots,0)$ is of type $\hat{s}_{1}$. In general a vertex of $D_{\mathcal{B}}$ is of type $\hat{s}_{i}$ if and only if the coordinate of this vertex has $i-1$ nonzero entries.
Let $\mathbb{D}_{\mathcal{B}}$ and $\mathbb{S}_{\mathcal{A}}$ be the Falk complex and spherical Deligne complex defined in Section 4.3. Then types of vertices in $D_{\mathcal{B}}$ and $S_{\mathcal{A}}$ pull back to type of vertices in $\mathbb{D}_{\mathcal{B}}$ and $\mathbb{S}_{\mathcal{A}}$. As in Section 5.2, we can view $\mathbb{D}_{\mathcal{B}}$ as a subcomplex of $\mathbb{S}_{\mathcal{A}}$ which is a connected component of the inverse image of $D_{\mathcal{B}}$ (viewed as a subcomplex of $S_{\mathcal{A}}$) under the map $\mathbb{S}_{\mathcal{A}}\to S_{\mathcal{A}}$.
Proposition A.1.
The vertex set of $\mathbb{D}_{\mathcal{B}}$, endowed with the relation $<$ induced from $s_{1}<s_{2}<\cdots<s_{n}$ as in Definition 2.15, is a poset satisfying all the assumptions of Theorem 2.18.
Proof.
Similar to the proof of Proposition 5.4, it suffices to verify Theorem 2.18 (3) for type $\hat{s}_{1}$ vertices and Theorem 2.18 (4) for type $\hat{s}_{n}$ vertices. If $x$ is of type $\hat{s}_{1}$, then $x$ maps to $\bar{x}=(0,0,\ldots,0)$ under $\mathbb{D}_{\mathcal{B}}\to D_{\mathcal{B}}$. Note that the local arrangement of $\mathcal{B}$ at $\bar{x}$ (as defined in Section 4.3) is an arrangement of type $B_{n-1}$. Then we are done by Theorem 2.37. If $x$ is of type $\hat{s}_{n}$, then up to symmetry, we can assume $x$ maps to $\bar{x}=(-1,-1,\ldots,-1)$ under $\mathbb{D}_{\mathcal{B}}\to D_{\mathcal{B}}$. The local arrangement of $\mathcal{B}$ at $\bar{x}$ is made of hyperplanes $\{y_{i}=-1\}_{i=1}^{n-1}$ and $\{y_{i}=y_{j}\}_{1\leq i\neq j\leq n-1}$, which is of type $A_{n-1}$ (up to a linear transformation). Then we can finish in the same way as the proof of Proposition 5.4, using Lemma 5.5, Lemma 4.10, and Theorem 2.37. ∎